Determine the length of the curve 𝑥 = 𝑦^2 /2 for 0 ≤ 𝑥 ≤ 1/2 . Assume 𝑦 positive.




1
Expert's answer
2022-01-10T17:50:37-0500

"L= \int \sqrt{1+(y')^2}dx\\
y=\sqrt{2x}\\
y'=\frac{1}{\sqrt{2x}}\\
L= \int_0^\frac{1}{2}\sqrt{1+\frac{1}{2x}}dx\\"

This can also be written as "L= \int_0^\frac{1}{2}\sqrt{\frac{1}{2x}+1}dx\\"

we apply linearity

"\frac{1}{\sqrt{2}}\int \sqrt{\frac{1}{x}+2}dx\\"

let u = "\sqrt{\frac{1}{x}+2}\\" , dx = -2"\sqrt{\frac{1}{x}+2x^2}du, dx=-2\int \frac{u^2}{(u^2-2)}du\\"

we integrate "\int \frac{u^2}{(u^2-2)}du = \int (\frac{u^2-2}{(u^2-2)^2}+\frac{2}{(u^2-2)^2})du=\\"

"\int (\frac{1}{(u^2-2)}+\frac{2}{(u^2-2)^2})du\\"

we split and integrate differently.

"\int (\frac{1}{(u^2-2)}du \implies \frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}}- \frac{Ln(u+\sqrt{2})}{2^{\frac{3}{2}}}"

also

"\int\frac{2}{(u^2-2)^2}du \implies \int (\frac{1}{(u^2-2)}du \implies \frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}}- \frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}} - \frac{1}{4(u+\sqrt{2})}- \frac{1}{4(u-\sqrt{2})}\\"

but substituting the solution of the integrations back, we get

"\frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}}- \frac{Ln(u+\sqrt{2})}{2^{\frac{3}{2}}}+ \frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}}- \frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}} - \frac{1}{4(u+\sqrt{2})}- \frac{1}{4(u-\sqrt{2})}"

recall, u= "\sqrt{\frac{1}{x}+2}\\" then we substitute u in terms of x back and introduce the lower and upper limits

"\frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}}- \frac{Ln(u+\sqrt{2})}{2^{\frac{3}{2}}}+ \frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}}- \frac{Ln(u-\sqrt{2})}{2^{\frac{3}{2}}} - \frac{1}{4(u+\sqrt{2})}- \frac{1}{4(u-\sqrt{2})}|_0^\frac{1}{2}"

We have.

"\frac{1}{\sqrt{2}}+\frac{Ln(\sqrt{2}+1)-Ln(\sqrt{2}-1)}{4}"